Concepts of trigonometric ratios

Grade:12th pass

1 Answers

Arpit Kumar
10 years ago
The trigonometric ratios are the basic building blocks of trigonometry so you need to know them before further studies. There are six trigonometric ratios also called trigonometric functions. 1>sine theta=p/h 2>cosine theta=b/h 3>tangent theta=sin theta/cos theta=p/b Their reciprocals in order. 1>cosecant theta=h/p 2>secant theta=h/b 3>cotangent theta=b/p Hope this helps